10x+7y=-10
-9x-7y=2
1
Expert's answer
2019-11-06T09:48:43-0500

"10x+7y=-10"

"-9x-7y=2"

Adding two equations

"x=-8"

Substituting for x into the first equation of the system

"10(-8)+7y=-10"

"-80+7y=-10"

"7y=-10+80"

"y=\frac {70} 7"

"y=10"

Answer: x=-8 and y=10